Suppose that the market demand in a perfectly competitive industry is given by Qd=70,000-5000P, and the market supply function is Qs=40,000+2500P given in dollors. the market equilibrium price will be
Answer by mathmate(429)


Problem:
In a competitive industry,
demand function: Qd=70000-5000P
supply function: Qs=40000+2500P
Find market equilibrium price.

Solution:
In a perfectly competitive market, demand = supply, so Qd=Qs.
Equate the two given equations and get
40000+2500P=70000-5000P
Transpose and solve for P
2500P+5000P=70000-40000
7500P=30000
P=4

Answer: The equilibrium price is $4.